An object falling from rest in air is subject not only to the gravitational force but also to air resistance. Assume that the air resistance is proportional to the velocity with constant of proportionality as

, and acts in a direction opposite to motion

Then velocity cannot exceed.
1
Expert's answer
2020-10-27T11:15:56-0400
"\frac{dv}{dt}=g-kv\\ln{(g-kv)}=C-kt\\g-kv>0\\v<\frac{k}{g}"